# Unlocking LLM Repair Capabilities in Low-Resource Programming Languages Through Cross-Language Translation and Multi-Agent Refinement

### v3

@@ 1. Introduction
By benefiting from training or fine-tuning (Hossain et al., 2024; Jin et al., 2023; Yuan et al., 2022; Luo et al., 2024) on diverse multilingual codebases, LLM-based repair approaches have achieved promising performance across different programming languages. However, these models exhibit substantial disparities in fixing bugs across languages, with significantly stronger repair capabilities in prevalent languages (2.0, 2024; Octoverse, 2022) such as Java and Python (Yuan et al., 2022; Joshi et al., 2023; Xia and Zhang, 2022) while struggling with others like Rust and Kotlin.
@@
It is important to note that programming problems often have equivalent implementations across different languages that serve the same underlying intentions and functions (Roziere et al., 2020; Tehrani et al., 2024), which suggests a natural opportunity: if a bug proves difficult to fix in one language, it might be beneficial to draw from the repair expertise available in another language. Intuitively, in other words, an LLM that exhibits superior repair capabilities in one language could be leveraged to assist with resolving issues in another. In addition, as the repair process unfolds new insights and historical experiences (Wen et al., 2019), these can be fed back into the repair cycle to guide future refinement. Inspired by these intuitions, we propose to leverage cross-language knowledge by translating the buggy code to other languages with iterative refinement using historical feedback.
Recent evidence (Ahmed and Devanbu, 2022) has revealed that LLMs trained in one programming language can sometimes even outperform those trained in the target language itself, suggesting valuable transfer effects across language boundaries. With rapid progress in code translation (Zhu et al., 2022; Roziere et al., 2020), the study of Pan et al. (Pan et al., 2024) demonstrates that LLMs can translate code better than non-LLM translation approaches for specific languages, demonstrating the potential of LLMs in code translation. Furthermore, benefited from advances in software engineering agents (Liu et al., 2024b; Yang et al., 2025), the integration of agent-driven (Xia and Zhang, 2024; Zhang et al., 2024c; Bouzenia et al., 2024; Hidvégi et al., 2024) and multi-agent (Lee et al., 2024; Zhang et al., 2024b) approaches can enhance program repair with historical feedback while autonomously navigating complex repair workflows throughout the iterative repair process. Therefore, we introduce the key hypothesis of our paper:
@@ 2.1. Code Translation
Code translation refers to the process of transforming source code from one programming language to another, mainly employed for code adaptation scenarios or migrating legacy codebases written in deprecated languages to more modern alternatives (Zhu et al., 2022). Traditional rule-based translation approaches are costly, requiring substantial manual effort from programmers with specialized expertise in both the source and target languages (Ramos et al., 2024). Existing translation tools typically specialize in specific language pairs (e.g., Java to C# (Chen et al., 2018)) and require significant domain-specific knowledge to develop and maintain (Bhatia et al., 2025). Meanwhile, unsupervised neural approaches, while promising, necessitate high-quality data for each language pair and computational resources for model training, making them also resource-intensive (Roziere et al., 2020). With LLMs demonstrating promising capability in translating code between programming languages (Pan et al., 2024), since program repair is the main objective of our paper rather than translation, we adopt LLM-based code translation for our approach to explore whether APR can be enhanced with code translation.
While a few recent research has adopted alternative terminology such as ”transpilation” (Ramos et al., 2024; Bhatia et al., 2025) or ”transcompilation” (Roziere et al., 2020), we find these terms largely interchangeable with ”translation” and thus use code translation throughout this paper, consistent with predominant usage in the literature.
@@ 2.2. LLM-Based Program Repair Agent
LLM-based agents employ LLMs as the cognitive core to perceive and act based on environmental feedback, working toward specific goals through four essential components: Planning, Memory, Perception, and Action (Liu et al., 2024b). The planning component generates plans with reasoning strategies tailored to the task. The memory component maintains a record of historical data, while the perception component processes environmental feedback to facilitate more effective planning. The action component executes concrete steps based on decisions made by the planning component. Building upon this basic framework, existing program repair agents such as ChatRepair (Xia and Zhang, 2024), AutoCodeRover (Zhang et al., 2024c), and RepairAgent (Bouzenia et al., 2024) typically follow a common pipeline consisting of four phases in an iterative refinement paradigm: (1) the agent first generates potential fixes for the buggy code, (2) these fixes undergo patch validation through compilation, execution, and automated testing tools autonomously, (3) the repair feedback from these validation steps is carefully analyzed to inform the next iteration of fixing, and (4) the process repeats until a solution passes all validation criteria or reaches a predetermined iteration limit. Moreover, multi-agent architectures further extend this paradigm by decomposing complex tasks into specialized modules. FixAgent (Lee et al., 2024) leverages multiple agents dedicated to autonomous bug fixing and bug localization. AgentCoder (Huang et al., 2023) integrates specialized agents responsible for code generation, test generation, and code execution. Similarly, ACFix (Zhang et al., 2024b) employs a configuration comprising a generator for proposing fixes and a validator tasked with ensuring their correctness. In our study, we leverage multiple LLM-based agents for program repair, code translation, and decision-making.
@@ 3. Illustrative Example
We illustrate how code translation can assist with repairing bugs that the LLM failed to fix with two real examples in our study. Figure 1 and Figure 2 present two examples with the original buggy code, failed fix with direct repair and the fixed code achieved through code translation, where the buggy code is translated to another programming language, attempted to be repaired in the target language, and then translated back to the original language.
@@
Figure 1 exhibits an example of a successful fix by translating the buggy Python code with a TIME LIMIT EXCEEDED error to C++. The original Python code iterates over every number in the interval $[a,b]$ and uses the condition bin(i).count(’0’)==2 to check for exactly one zero in the binary representation. This approach not only becomes inefficient for large intervals, which lead to a time exceeding error, but also suffers from a logical imprecision since Python’s bin() function adds prefixes to the result with ”0b”, distorting the intended count. Although the failed fix by direct repair attempted to mitigate this issue by stripping off the prefix with bin(i)[2:], the brute-force iteration over the entire range remains a performance bottleneck for time efficiency. In contrast, translating the code to C++ prompts a complete re-implementation that leverages low-level operations, particularly through bit shifts (<<much-less-than$<$ <) and direct arithmetic. This translation restricts the computation to iterating over possible bit lengths (from 2 to 63) and candidate positions for the single zero, rather than every single number, thereby dramatically reducing the search space. Technically, C++’s built-in facilities such as bitset and its emphasis on low-level efficiency not only expose latent inefficiencies in the original language but also foster an algorithmic design that is both conceptually elegant and computationally superior.
@@
Figure 2 demonstrates a successful fix by translating a C++ bug with MEMORY LIMIT EXCEEDED error to Rust. The original C++ code fails due to an oversized allocation (array size 1000) for factorial computations and unsafe concurrent collection modification. Direct C++ repair only reduces the factorial table size to 20 without addressing the iteration issue. When translated to Rust, the language’s strict safety guarantees force a re-examination of these problems as they violate Rust’s rules. After translating back to C++, the solution incorporates multiple improvements: reduced table size, safe iteration logic, and enhanced type safety through $int64\_t$ adoption. While the algorithm remains similar, Rust’s language-specific features regarding memory safety help identify and fix memory and logical flaws that are less apparent in C++.
@@ 5.1. Research Questions
5.2. Model, Benchmark, and Metrics
@@ 5.2.1. Model
We use the open-source state-of-the-art DeepSeek-V3 (Liu et al., 2024a) model as the LLM in the analyzer and translator, for both decision-making and code translation. We also leverage the same LLM for the repairer to align the language scope with the translator instead of adopting various APR approaches specialized for different languages of bugs. We choose DeepSeek-V3 because it is arguably among the best for coding tasks in the literature.
@@ 5.2.2. Benchmark
To comprehensively evaluate LANTERN, we employ xCodeEval (Khan et al., 2024), a state-of-the-art multilingual benchmark covering 7 code-related tasks, including program repair with 11 different programming languages. xCodeEval is constructed from 25 million openly available samples on Codeforces (Codeforces, 2024), which cover 7,514 unique problems with fine-grained difficulty ratings ranging from 800 to 3,500 points (according to Codeforces), enabling evaluation across varying levels of code complexity. We use one of the subsets provided by xCodeEval, Compact Set, which is dedicated to academic research. The Compact Set comprises a total of 5,068 bugs across 11 languages, which are presented in Table 1. Moreover, each problem in xCodeEval is accompanied by comprehensive unit tests averaging 50 tests per problem, substantially exceeding some widely used alternatives, such as Humaneval (Chen et al., 2021), which only provides seven test cases per problem on average and covers fewer programming languages. In addition, xCodeEval’s execution engine, ExecEval, provides dedicated runtime environments for all necessary compilers and interpreters, categorizing bug executions into six outcomes: compilation error, runtime error, memory limit exceeded, time limit exceeded, wrong answer, and passed.
@@
{tblr}
width = colspec = Q[63]Q[63]Q[75]Q[63]Q[71]Q[63]Q[94]Q[71]Q[112]Q[85]Q[75]Q[81], hlines, C & C# C++ Go Java JS Kotlin PHP Python Ruby Rust Total
@@ 5.2.3. Metrics
Pass@k is a commonly used metric that measures the likelihood that at least one candidate sample among a set of $k$ generated samples fixes a bug (Chen et al., 2021; Zheng et al., 2023), which is calculated as follows:
@@
where $n$ is the total number of samples generated for each problem, and $c$ is the number of correct samples (with $k\leq n$ and $c\leq n$). According to the finding of Noller et al. (Noller et al., 2022), which indicates that developers are generally willing to review a maximum of approximately 10 patches, we use $Pass@10$ with a total of $n=20$ samples generated for each bug.
@@
Precision@k quantifies the proportion of valid iterations in the top $k$ iterations. For the iterations in which the bugs of a given language $q$ are being repaired, with $Q$ being the set of all programming languages, let the ranked list of $k$ candidate iterations be denoted as $[l_{1},l_{2},\ldots,l_{k}]$. The $Precision@k$ metric is computed as:
(2) $\small Precision@k$ $=\frac{1}{|Q|}\sum_{q\in Q}Precision@k(q)$
(3) $=\frac{1}{|Q|}\sum_{q\in Q}\frac{\sum_{j=1}^{k}\textbf{rel}(q,l_{j})}{k}.$
where $\textbf{rel}(q,l_{j})$ is a binary indicator that is $1$ if valid fixes exist in the candidate iteration $l_{j}$ of language $q$, and 00 otherwise.
Mean Average Precision at k (MAP@k) evaluates the overall quality of the rankings of the target languages. Firstly, $AP@k$ across $k$ iterations of repair for bugs of language $q$ is defined as:
(4) $\small AP@k(q)=\frac{1}{\min(|R_{q}|,k)}\sum_{j=1}^{k}Precision@j(q)\cdot\textbf{rel}(q,l_{j}),$
where $|R_{q}|$ is the total number of valid iterations for the repair of bugs in language $q$. The $MAP@k$ is then calculated as:
(5) $\small MAP@k=\frac{1}{|Q|}\sum_{q\in Q}AP@k(q).$
Normalized discounted cumulative gain at k (NDCG@k) measures the ranking quality by giving higher weights to valid iterations that appear earlier. Firstly, the $DCG@k$ for language $q$ is computed as:
(6) $\small DCG@k(q)=\sum_{j=1}^{k}\frac{\textbf{rel}(q,l_{j})}{\log_{2}{(j+1)}},$
and the ideal $DCG@k$ ($IDCG@k$) is the maximum possible $DCG@k$ obtainable by a perfect ranking where all valid iterations occur at the top-$k$ positions:
(7) $\small IDCG@k(q)=\sum_{j=1}^{\min(|R_{q}|,k)}\frac{1}{\log_{2}{(j+1)}},$
The $NDCG@k$ is then given by:
(8) $NDCG@k$ $=\frac{1}{|Q|}\sum_{q\in Q}NDCG@k(q)$
(9) $=\frac{1}{|Q|}\sum_{q\in Q}\frac{DCG@k(q)}{IDCG@k(q)}.$
Recall@k measures the proportion of valid iterations among the top $k$ iterations in the number of all relevant iterations, which is defined as:
(10) $Recall@k$ $=\frac{1}{|Q|}\sum_{q\in Q}Recall@k(q)$
(11) $=\frac{\sum_{j=1}^{k}\textbf{rel}(q,l_{j})}{|R_{q}|}.$
@@
In addition, we employ the Mann-Whitney-Wilcoxon (MWW) (Mann and Whitney, 1947; Wilcoxon, 1992) test to determine the statistical significance of differences in performance metrics across various strategies, with Cliff’s Delta (Cliff, 1993) quantifying the magnitude of these differences. The Cliff’s Delta is confined to the range $[-1,1]$, with 0 indicating no difference and values closer to -1 or 1 signifying a larger effect. Following the study of Yang et al. (Yang et al., 2024), we interpret the absolute value of Cliff’s Delta as follows: (1) negligible effect: $[0,0.11)$; (2)small effect: $[0.11,0.28)$; (3) medium effect: $[0.28,0.43)$; and (4) large effect: $[0.43,1]$.
5.3. RQ1: Overall Effectiveness
5.3.1. Experimental Design

### v4

# Unlocking LLM Repair Capabilities Through Cross-Language Translation and Multi-Agent Refinement

@@ 1. Introduction
By benefiting from training or fine-tuning (Hossain et al., 2024; Jin et al., 2023; Yuan et al., 2022; Luo et al., 2024) on diverse multilingual codebases, LLM-based repair approaches have achieved promising performance. However, these models exhibit substantial disparities in fixing bugs across languages, with significantly stronger repair capabilities in prevalent languages (2.0, 2024; Octoverse, 2022) such as Java and Python (Yuan et al., 2022; Joshi et al., 2023; Xia and Zhang, 2022) while struggling with others like Rust and Kotlin.
@@
Programming problems often have equivalent implementations across different languages that serve the same underlying intentions and functions (Roziere et al., 2020; Tehrani et al., 2024), which suggests a natural opportunity: if a bug proves difficult to fix in one language, it might be beneficial to draw from the repair expertise available in another language. Intuitively, in other words, an LLM that exhibits superior repair capabilities in one language could be leveraged to assist with resolving issues in another. In addition, as the repair process unfolds new insights and historical experiences (Wen et al., 2019), these can be fed back into the repair cycle to guide future refinement. Inspired by these intuitions, we propose to leverage cross-language knowledge by translating the buggy code to other languages with iterative refinement using historical feedback.
LLMs trained in one language can sometimes even outperform those trained in the target language itself (Ahmed and Devanbu, 2022), suggesting valuable transfer effects across language boundaries. With rapid progress in code translation (Zhu et al., 2022; Roziere et al., 2020), the study of Pan et al. (Pan et al., 2024) demonstrates that LLMs can translate code better than non-LLM translation approaches for specific languages, demonstrating the potential of LLMs in code translation. Furthermore, benefited from advances in software engineering agents (Liu et al., 2024b; Yang et al., 2025), the integration of agent-driven (Xia and Zhang, 2024; Zhang et al., 2024c; Bouzenia et al., 2024; Hidvégi et al., 2024) and multi-agent (Lee et al., 2024; Zhang et al., 2024b) approaches can enhance program repair with historical feedback while autonomously navigating complex repair workflows throughout the iterative repair process. Therefore, we introduce the key hypothesis:
@@ 2.1. Code Translation
Code translation refers to the process of transforming source code from one programming language to another, mainly employed for code adaptation scenarios or migrating legacy codebases written in deprecated languages to more modern alternatives (Zhu et al., 2022). Traditional rule-based translation approaches are costly, requiring substantial manual effort from programmers with specialized expertise in both the source and target languages (Ramos et al., 2024). Existing translation tools typically specialize in specific language pairs (e.g., Java to C# (Chen et al., 2018)) and require significant domain-specific knowledge to develop and maintain (Bhatia et al., 2025). Meanwhile, unsupervised neural approaches, while promising, necessitate high-quality data for each language pair and computational resources for model training, making them also resource-intensive (Roziere et al., 2020).
LLMs has shown promising capability in translating code between programming languages (Pan et al., 2024). However, recent research highlights a significant gap between translating isolated code snippets and migrating entire software repositories, which involves complex dependencies, build configurations, and project-specific architectures (Wang et al., 2024; Zhang et al., 2025b; Ou et al., 2024; Yan et al., 2023). To tackle this, the field has shifted towards repository-level translation, developing new benchmarks and neuro-symbolic methods that combine program analysis with LLMs to manage complexity (Ibrahimzada et al., 2025; Zhang et al., 2025a). These studies show that while fully automated, correct translation of large projects remains a challenge, the generated code can significantly reduce manual effort. Since program repair is the main objective of our paper rather than translation, we adopt LLM-based code translation for our approach to explore whether APR can be enhanced with code translation.
@@ 2.2. LLM-Based Program Repair Agent
LLM-based agents employ LLMs as the cognitive core to perceive and act based on environmental feedback, working toward specific goals through four essential components: Planning, Memory, Perception, and Action (Liu et al., 2024b). The planning component generates plans with reasoning strategies tailored to the task. The memory component maintains a record of historical data, while the perception component processes environmental feedback to facilitate more effective planning. The action component executes concrete steps based on decisions made by the planning component. Building upon this, existing program repair agents such as ChatRepair (Xia and Zhang, 2024), AutoCodeRover (Zhang et al., 2024c), and RepairAgent (Bouzenia et al., 2024) typically follow a pipeline consisting of four steps in an iterative paradigm: (1) the agent first generates potential fixes for the buggy code, (2) these fixes undergo patch validation through compilation, execution, and automated testing autonomously, (3) repair feedback from validation step is carefully analyzed to inform the next iteration of fixing, and (4) the process repeats until a solution passes all validation criteria or reaches a predetermined iteration limit. Moreover, multi-agent architectures further extend this paradigm by decomposing complex tasks into specialized modules. FixAgent (Lee et al., 2024) leverages multiple agents dedicated to autonomous bug fixing and bug localization. AgentCoder (Huang et al., 2023) integrates specialized agents responsible for code generation, test generation, and code execution. Similarly, ACFix (Zhang et al., 2024b) employs a configuration comprising a generator for proposing fixes and a validator tasked with ensuring their correctness. In our study, we leverage multiple LLM-based agents for program repair, code translation, and decision-making.
@@ 3. Illustrative Example
We illustrate how code translation can assist with repairing bugs that the LLM failed to fix with two real examples in our study. Figure 1 and Figure 2 present two examples demonstrating the the efficiency of program repair with code translation. The examples show (1) the original buggy code, (2) the failed fix from direct repair, (3) the translated buggy code in a new programming language, (4) the fixed code in the target language, and (5) the back-translated fixed code. This detailed breakdown shows how the buggy code is first translated, then repaired in the target language, and finally translated back to its original language.
@@
Figure 1 exhibits an example of a successful fix by translating the buggy Python code with a TIME LIMIT EXCEEDED error to C++. The original Python code iterates over every number in the interval $[a,b]$ and uses the condition bin(i).count(’0’)==2 to check for exactly one zero in the binary representation. This approach not only becomes inefficient for large intervals, which lead to a time exceeding error, but also suffers from a logical imprecision since Python’s bin() function adds prefixes to the result with ”0b”, distorting the intended count. Although the failed fix by direct repair attempted to mitigate this issue by stripping off the prefix with bin(i)[2:], the brute-force iteration over the entire range remains a performance bottleneck for time efficiency. When translated to C++, the buggy logic is preserved. The resulting C++ code still uses a brute-force loop from a to b and checks the property of each number, meaning it would also fail with a TIME LIMIT EXCEEDED error. When prompted to fix the inefficient C++ code, the LLM devises a complete re-implementation that leverages low-level operations, particularly through bit shifts (<<<<) and direct arithmetic. The fixed code in C++ restricts the computation to iterating over possible bit lengths (from 2 to 63) and candidate positions for the single zero, rather than every single number, thereby dramatically reducing the search space. Finally, this optimized algorithm is then successfully translated back into the back-translated fixed code in Python, resolving the original time exceeding error. Technically, C++’s built-in facilities such as bitset and its emphasis on low-level efficiency not only expose latent inefficiencies in the original language but also foster an algorithmic design that is both conceptually elegant and computationally superior.
Figure 2 presents a successful fix by translating a C++ bug with MEMORY LIMIT EXCEEDED error to Rust. The original C++ code fails due to oversized factorial allocation (array size 1000) and unsafe concurrent collection modification. Direct C++ repair only reduces the table size to 20, leaving the iteration issue unresolved. When this code is translated to Rust, the language’s strict safety and ownership rules immediately expose the latent bug. The translated code, which attempts to remove an item from a collection while iterating over it, would be rejected by the Rust compiler due to borrow-checking rules that prevent such unsafe operations. Translating to Rust forces a re-examination of the flawed logic with the language’s strict safety guarantees. After translating back to C++, the solution incorporates multiple improvements: reduced table size, safe iteration logic, and enhanced type safety through int64_tint64\_t adoption. While the algorithm remains similar, Rust’s language-specific features regarding memory safety help identify and fix memory and logical flaws that are less apparent in C++.
@@ 5.1. Research Questions
  • RQ1.3 (Approach Comparison): How does the repair performance of LANTERN compare against other state-of-the-art (SOTA) program repair approaches?
@@
RQ4 (Generalizability):
  • RQ4.1 (Real-World Generalizability): How effectively does LANTERN generalize to real-world APR benchmarks?
  • RQ4.2 (Model Generalizability): Can LANTERN generalize to other LLMs and to what extent does LANTERN’s effectiveness generalize across different LLMs?
5.2. Experiment Setup
@@ 5.2.1. Model
For primary experiments from RQ1 to RQ3, we use the open-source state-of-the-art DeepSeek-V3 (Liu et al., 2024a) model as the LLM in the analyzer and translator, for both decision-making and code translation. We also leverage the same LLM for the repairer to align the language scope with the translator instead of adopting various APR approaches specialized for different languages of bugs. We choose DeepSeek-V3 because it is arguably among the best for coding tasks in the literature.
To validate the generalizability of our approach, we include two additional prominent LLMs in RQ4. We selected one closed-source model, Claude 3.5 Sonnet (Anthropic, 2024) and another powerful open-source model, Qwen2.5-72B-Instruct (Team, 2024; Yang et al., 2024b), as a SOTA open-source alternative. This selection allows us to assess whether our approach’s benefits are tied to a specific model architecture or are more broadly applicable across different foundational models.
@@ 5.2.2. Benchmark
We employ three distinct and challenging benchmarks. First, to assess our approach from RQ1 to RQ3, we use xCodeEval (Khan et al., 2024), a state-of-the-art multilingual benchmark covering 7 code-related tasks, including program repair with 11 different programming languages. xCodeEval is constructed from 25 million openly available samples on Codeforces (Codeforces, 2024), which cover 7,514 unique problems with fine-grained difficulty ratings ranging from 800 to 3,500 points (according to Codeforces), enabling evaluation across varying levels of code complexity. We use one of the subsets provided by xCodeEval, Compact Set, which is dedicated to academic research. The Compact Set comprises a total of 5,068 bugs across 11 languages, which are presented in Table 1. Moreover, each problem in xCodeEval is accompanied by comprehensive unit tests averaging 50 tests per problem, substantially exceeding some widely used alternatives, such as Humaneval (Chen et al., 2021), which only provides seven test cases per problem on average and covers fewer programming languages. In addition, xCodeEval’s execution engine, ExecEval, provides dedicated runtime environments for all necessary compilers and interpreters, categorizing bug executions into six outcomes: compilation error, runtime error, memory limit exceeded, time limit exceeded, wrong answer, and passed.
Second, we extend our evaluation scenario to real-world projects in RQ4. We incorporate Defects4J (Just et al., 2014), a widely-used benchmark for Java program repair. Defects4J 1.2 contains 391 (removing 4 deprecated bugs) real-world bugs from 6 large-scale Java projects and Defects4J 2.0 contains 438 bugs from 9 additional projects. In addition to Defects4J, to assess performance on a more complex repository-level APR setting, we utilize the SWE-Bench (Jimenez et al., 2023) benchmark. SWE-Bench presents a significant challenge by tasking models with resolving real-world GitHub issues from popular Python repositories such as Django and scikit-learn. Specifically, we use the light-weight subset, SWE-Bench Lite that comprises 300 instances across 11 repositories from the full benchmark. The benchmark is notably difficult, with SOTA models achieving relatively low resolve rates compared to other benchmarks, making it an excellent choice for evaluating the real-world generalizability of our approach.
@@
C C# C++ GoGo Java JS Kotlin PHPPHP Python Ruby Rust Total
@@ 5.2.3. Metrics
Pass@k measures the likelihood that at least one of the top $k$ generated patches fixes a bug (Chen et al., 2021; Zheng et al., 2023):
@@
with $n$ patches and $c$ correct samples, where $k\leq n$ and $c\leq n$. Since developers are generally willing to review a maximum of approximately 10 patches (Noller et al., 2022), we use $Pass@10$ under $n=20$.
@@
Precision@k quantifies the proportion of valid iterations in the top $k$ iterations. For the iterations in which the bugs of a given language $q$ are being repaired, with $Q$ being the set of all programming languages, let the ranked list of $k$ candidate iterations be denoted as $l_{1},l_{2},\ldots,l_{k}]$.
Mean Average Precision at k (MAP@k) evaluates the overall quality of the rankings of target languages, which is calculated as:
(2) $\small MAP@k=\frac{1}{|Q|}\sum_{q\in Q}\frac{1}{\min(|R_{q}|,k)}\sum_{j=1}^{k}Precision@j(q)\cdot\textbf{rel}(q,l_{j}).$
Normalized discounted cumulative gain at k (NDCG@k) measures the ranking quality by giving higher weights to valid iterations that appear earlier.
Recall@k measures the proportion of valid iterations among the top $k$ iterations in the number of all relevant iterations.